# HEIGHT: Heterogeneous Interaction Graph Transformer for Robot Navigation in Crowded and Constrained Environments

### v3

@@ Abstract
footnotetext: Manuscript received July, 10, 2023; Revised August 8, 2025 and October 12, 2025; Accepted November 13, 2025; Date of publication December 15, 2025; Date of current version December 15, 2025. This article was recommended for publication by Associate Editor and Editor Chenguang Yang upon evaluation of the reviewers’ comments. (Corresponding author: Shuijing Liu.)
@@ I Introduction
While a large body of research on robot crowd navigation studies open-world navigation and ignores static obstacles [chen2019crowd, liu2023intention, wang2023navistar], some previous works have explored various approaches for crowd navigation in constrained spaces [fox1997dynamic, van2008reciprocal, perez2021robot]. However, these works typically have one of two limitations: (1) They do not differentiate between dynamic and static obstacles, and thus the robot has difficulties taking adaptive strategies to avoid collisions [fox1997dynamic, chen2019relational, perez2021robot, liu2020decentralized]; (2) The representations of scene, especially static obstacles, are inefficient or redundant as inputs to the navigation algorithm. To address these limitations, we ask the following research question: How can a robot navigation policy represent and reason about diverse interactions in crowded and constrained environments to adaptively avoid collisions?
To answer this question, we propose a framework that leverages the heterogeneity of interactions in crowded and constrained scenarios in both scene representation and navigation policy design. First, we use light weight perception modules to extract separate human and obstacle representations that are essential for navigation, such as relative positions w.r.t. the robot. The split representations are processed and fed separately into the reinforcement learning (RL)-based navigation pipeline. Then, inspired by recent breakthroughs in spatio-temporal (st) networks for crowd navigation [liu2020decentralized, wang2023navistar, liu2023intention, chen2024socially], we decompose the scenario into a heterogeneous st-graph with different types of edges to represent different types of interactions among the robot, detected humans, and observed obstacles, as shown in the colored arrows in Fig. 1. Finally, we convert the heterogeneous st-graph into a HEterogeneous Interaction GrapH Transformer (HEIGHT), a robot policy network consisting of different modules to parameterize the various spatio-temporal interactions. Specifically, we use two separate attention networks to address the different effects of robot-human (RH) and human-human (HH) interactions. The attention networks enable the robot to pay more attention to the important interactions, leading to a low collision-rate even as the number of humans increases.
@@ II-A Model-based methods
Robot navigation in human crowds is particularly challenging and has been studied for decades [fox1997dynamic, hoy2015algorithms, savkin2014seeking, mavrogiannis2021core]. Model-based approaches have explored various mathematical models to optimize robot actions [van2011reciprocal, van2008reciprocal, fox1997dynamic, helbing1995social, huber2022avoiding, mavrogiannis2023winding]. As an early example, ROS navigation stack [ros_navigation_stack] uses a cost map for global planning and dynamic window approach (DWA) for local planning [fox1997dynamic]. DWA searches for the optimal velocity that brings the robot to the goal with the maximum clearance from any obstacle while obeying the dynamics of the robot. By treating humans as obstacles, DWA exhibits myopic behaviors such as oscillatory paths in dynamic environments [dobrevski2024dynamic]. As a step forward, optimal reciprocal collision avoidance (ORCA) and social force account for the velocities of agents. ORCA models other agents as velocity obstacles and assumes that agents avoid each other under the reciprocal rule [van2011reciprocal, van2008reciprocal]. SF models the interactions between the robot and other agents using attractive and repulsive forces [helbing1995social]. However, the hyperparameters of the model-based approaches are sensitive to crowd behaviors and thus need to be tuned carefully to ensure good performance [long2018towards, dobrevski2024dynamic]. In addition, model-based methods are prone to failures, such as the freezing problem, if the assumptions such as the reciprocal rule are broken [trautman2010unfreezing, liu2024relaxing]. In contrast, while our method also models these interactions, we learn the hyperparameters of the model from trial and error with minimal assumptions on human behaviors posed by model-based methods.
@@ II-B Learning-based methods
Learning-based approaches have been widely used for navigation in dynamic environments to reduce hyperparameter tuning efforts and the number of assumptions. One example is supervised learning from expert demonstrations of desired behaviors, where the expert ranges from model-based policies [long2017deep, tai2018socially, xie2021towards], human teleoperators in simulators [pokle2019deep], to real pedestrians [karnan2022scand, chandra2024towards]. Supervised learning does not require explorations of the state and action spaces of the robot, yet the performance of learned policy is limited by the quality of expert demonstrations.
Another line of work takes advantage of crowd simulators and learns policies with RL. Through trial and error, RL has the potential to learn robot behaviors that outperform model-based approaches and expert demonstrations [xie2023drlvo]. For example, Deep V-Learning first uses supervised learning with ORCA as the expert and then uses RL to learn a value function for path planning [chen2017decentralized, everett2018motion, chen2019crowd, chen2020robot_gaze, yang2023st2]. However, Deep V-Learning assumes that state transitions of all agents are known without uncertainty. In addition, since the networks are pre-trained with supervised learning, they share the same disadvantages with OCRA, which are hard to correct by RL [liu2020decentralized]. To address these problems, more recent efforts leverage model-free RL without supervised learning or assumptions on state transitions [liu2020decentralized, mun2023occlusion, liu2024relaxing, yang2023st2]. Since the state transition probability of humans is uncertain, directly learning a policy network without explicitly modeling state transitions is more suitable for navigation [sathyamoorthy2020densecavoid]. However, the aforementioned RL works have at least one of the following two problems: (1) They focus on navigation in open spaces and isolate agents from static constraints, posing difficulties when deploying robots in the real-world; (2) They ignore all or part of interactions among entities in the scene, which are important for robot navigation in dense crowds and highly constrained environments. We discuss how prior works address the above two problems in Sec. II-C and Sec. II-D respectively.
@@ II-C Scene representation for constrained environments
To deal with both static constraints and humans, some methods such as DWA [fox1997dynamic] and DS-RNN [liu2020decentralized, chen2024decentralized] use groups of small circles to represent the contours of both static obstacles and humans. While straightforward, groups of circle representation has the following problems: 1. Humans and obstacles are treated in the same way but have different effects on robot decision making. This design choice leads to suboptimal policy as we will show in Sec. V and VI. 2. Since the length of input vectors is proportional to the number of obstacles and humans, the hyperparameters of navigation algorithms easily overfit to a specific environment setting, causing performance degradation when the densities of obstacles and humans change.
Other learning-based approaches use raw sensor images or point clouds to represent the whole environment [perez2021robot, dugas2022navdreams, pokle2019deep, zheng2022hierarchical, sathyamoorthy2020densecavoid]. These end-to-end (e2e) pipelines have made promising progress in simulation. However, generalizing these e2e methods to real-world scenarios requires high-fidelity simulators [tsoi2022sean2, puig2024habitat, Kstner2024Arena3A] and/or datasets [martin2021jrdb, karnan2022scand, hirose2023sacson]. Since accurate pedestrian simulation is technically difficult and expensive [Kstner2024Arena3A], learning a deployable e2e policy in dense and interactive crowds remains an open challenge. To circumvent this problem, other works leverage processed state information such as occupancy maps [xie2023drlvo, yang2023rmrl]. However, compared to human circles or point clouds, occupancy maps expand the state space, which reduces training efficiency. In addition, the discretization of agent positions limits the ability of occupancy maps to capture slow agent movements. For this reason, we develop a structured input representation, consisting of human circles and processed obstacle point cloud, that only includes the most essential information for navigation. This input representation leads to strong performance in both simulation and real-world while only requiring a low-fidelity simulator.
@@ II-D Graph attention networks for multi-agent interactions
In recent years, attention mechanisms have demonstrated success in various fields [vaswani2017attention, velickovic2018graph, vemula2018social]. Vaswani et al. propose a transformer with self-attention mechanism that achieves state-of-the-art performance in machine translation [vaswani2017attention]. Later, graph attention networks show the effectiveness of attention on learning relationships of data with graphical structures such as citation networks [velickovic2018graph]. Inspired by these works, researchers in trajectory prediction and crowd navigation have found that attention networks are also well-suited to capture interactions amongst agents and entities, which contain essential information for multi-agent tasks [vemula2018social, huang2019stgat, chen2019crowd, liu2020decentralized, hasan2022meta]. For each agent, these works compute attention scores for all neighboring agents. Due to the permutation invariance property, attention scores better capture the importance of pair-wise relationships than encoding the states of all agents with a concatenation operation or an recurrent network [chen2017decentralized, everett2018motion, chen2024decentralized].
@@
More specifically, a line of works uses a robot-human attention network to determine the relative importance of each human to the robot [chen2019crowd, leurent2019social, liu2020decentralized, liu2021learning]. However, interactions among humans, which can also influence the robot, are not explicitly modeled. To this end, another line of works use a homogenous graph network to include both RH and HH interactions [chen2019relational, yang2023st2, liu2024sample]. However, since these works feed RH and HH features to a single attention network, the resulting robot policy has difficulty reasoning the specificity of each type of feature, which limits the robot’s ability to adapt to different interactions, as demonstrated in [chen2024socially] and our experiments in Sec. V. To deal with this issue, some recent works use multiple attention networks to encode RH and HH features separately, leading to improved navigation performance in dense and interactive crowds [liu2023intention, wang2023navistar, yang2023st2]. However, these works only deal with open-world social navigation and thus ignore the interactions between agents and obstacles, which are common in real-world navigation scenarios.
@@
To handle obstacles, Liu et al. [liu2023graph] and Chen et al. [chen2024socially] treat the humans, the robot, and static objects as nodes in graph neural networks to model the spatio-temporal interactions among all observed entities. Different from Chen et al. [chen2024socially] where a robot needs to reach an object such as a chair, the robot in this work only needs to avoid all objects. Thus, Chen et al. [chen2024socially] represent static obstacles as a semantic map with a class label for each object, whereas we represent all observed static obstacles as a point cloud. In Liu et al. [liu2023graph], representing each obstacle as a geometric shape requires further map processing, such as obstacle segmentation. Thus, we treat the point cloud consisting of all obstacles as a single node in our st-graph, which reduces both the processing requirement and the size of the graph.
@@ III-B Scene representation
To aid policy learning with the complicated state space involving multiple agents and entities, we develop a structured representation that splits human and obstacle states to account for their different effects on robot decision making. In Fig. 3, at each timestep $t$, we split a scene into a human representation ${h}^{t}_{1},...,{h}^{t}_{n}$, and an obstacle representation ${o}^{t}$. Each human has its own state since it can move independently from one another, whereas all obstacles are mixed together since they are all static. In human representation (Fig. 3(b)), the position and velocity of each human are detected using off-the-shelf human detectors [Jia2020DRSPAAM, yolov3, deepsort]. Tracking each human is not required, further simplifying perception requirements. By representing each human as a low-dimentional state vector, we abstract away detailed information such as gaits and appearance, which are difficult to simulate accurately [tsoi2022sean2, mavrogiannis2023core].
@@ IV-A Heterogeneous Spatio-Temporal Graph
The subtle yet highly dynamic interactions among agents and entities are important factors that makes crowd navigation difficult. To model these interactions in a structured manner, we formulate the navigation scenario as a heterogeneous st-graph. In Fig. 4a, at each timestep $t$, our heterogeneous st-graph $\mathcal{G}^{t}=(\mathcal{V}^{t},\mathcal{E}^{t})$ consists of a set of nodes $\mathcal{V}^{t}$ and a set of edges $\mathcal{E}^{t}$. The nodes include the detected humans $h_{1}^{t},...,h_{n}^{t}$ and the robot $w^{t}$. In addition, an obstacle node $o^{t}$ represents the point cloud of all obstacles as a whole. At each timestep $t$, the spatial edges that connect different nodes denote the spatial interactions among nodes. Different spatial interactions have different effects on robot decision-making. Specifically, since we have control of the robot but not the humans, RH interactions have direct effects while HH interactions have indirect effects on the robot actions. As an example of indirect effects, if human A aggressively forces human B to turn toward the robot’s front, the robot has to respond as a result of the interaction between A and B. Additionally, since the agents are dynamic but the obstacles are static, interactions among agents are mutual while the influence of static obstacles on agents is one-way. Thus, we categorize the spatial edges into three types: HH edges (blue in Fig. 4), obstacle-agent (OA) edges (orange), and RH edges (red). The three types of edges allow us to factorize the spatial interactions into HH, OA, and RH functions. Each function is parameterized by a neural network that has learnable parameters. Compared to previous works that ignore some edges [chen2019crowd, liu2020decentralized, liu2023intention], our method allows the robot to reason about all observed spatial interactions that exist in constrained and crowded environments.
@@
To reduce the number of parameters, the same type of edges in Fig. 4(a) share the same function parameters. This parameter sharing is important for the scalability of our graph because the number of parameters is kept constant when the number of human changes [jain2016structural].
@@ IV-B1 Attention among agents
The attention modules assign weights to all edges that connect to a robot or a human node, allowing the node to pay attention to important edges or interactions. The two attention networks are similar to the scaled dot-product attention with a padding mask [vaswani2017attention], which computes the attention score using a query $Q$ and a key $K$, and applies the normalized score to a value $V$, which results in a weighted value $v$.
@@
where $c_{1},...,c_{n}$ are constants that reflect the effect of the scaling factor $d$ and the softmax function. Thus, the value of the $n$-th missing human $v_{n}$ is eliminated by the mask $M^{t}$ and thus does not affect the resulting weighted human embedding $v_{HH}^{t}$. The mask that indicates the visibility of each human prevents attention to undetected humans, which is common in crowd navigation due to the partial observability caused by the limited robot sensor range, occlusions, imperfect human detectors, etc [mun2023occlusion]. Additionally, the mask provides unbiased gradients to the networks, which stabilizes and accelerates the training [ma2020reinforcement, liu2023intention].

### v4

@@ Abstract
footnotetext: Manuscript received July, 10, 2023; Revised August 8, 2025 and October 12, 2025; Accepted November 13, 2025; Date of publication December 15, 2025; Date of current version December 15, 2025. This article was recommended for publication by Associate Editor and Editor Chenguang Yang upon evaluation of the reviewers’ comments.footnotetext: Corresponding author: Shuijing Liu (shuijing.liu@utexas.edu).footnotetext: Digital Object Identifier (DOI): 10.1109/TASE.2025.3646588.
@@ I Introduction
While a large body of research on robot crowd navigation studies open-world navigation and ignores static obstacles [7, 35, 65], some previous works have explored various approaches for crowd navigation in constrained spaces [17, 60, 48]. However, these works typically have one of two limitations: (1) They do not differentiate between dynamic and static obstacles, and thus the robot has difficulties taking adaptive strategies to avoid collisions [17, 6, 48, 36]; (2) The representations of scene, especially static obstacles, are inefficient or redundant as inputs to the navigation algorithm. To address these limitations, we ask the following research question: How can a robot navigation policy represent and reason about diverse interactions in crowded and constrained environments to adaptively avoid collisions?
To answer this question, we propose a framework that leverages the heterogeneity of interactions in crowded and constrained scenarios in both scene representation and navigation policy design. First, we use light weight perception modules to extract separate human and obstacle representations that are essential for navigation, such as relative positions w.r.t. the robot. The split representations are processed and fed separately into the reinforcement learning (RL)-based navigation pipeline. Then, inspired by recent breakthroughs in spatio-temporal (st) networks for crowd navigation [36, 65, 35, 5], we decompose the scenario into a heterogeneous st-graph with different types of edges to represent different types of interactions among the robot, detected humans, and observed obstacles, as shown in the colored arrows in Fig. 1. Finally, we convert the heterogeneous st-graph into a HEterogeneous Interaction GrapH Transformer (HEIGHT), a robot policy network consisting of different modules to parameterize the various spatio-temporal interactions. Specifically, we use two separate attention networks to address the different effects of robot-human (RH) and human-human (HH) interactions. The attention networks enable the robot to pay more attention to the important interactions, leading to a low collision-rate even as the number of humans increases.
@@ II-A Model-based methods
Robot navigation in human crowds is particularly challenging and has been studied for decades [17, 22, 54, 44]. Model-based approaches have explored various mathematical models to optimize robot actions [59, 60, 17, 20, 27, 43]. As an early example, ROS navigation stack [47] uses a cost map for global planning and dynamic window approach (DWA) for local planning [17]. DWA searches for the optimal velocity that brings the robot to the goal with the maximum clearance from any obstacle while obeying the dynamics of the robot. By treating humans as obstacles, DWA exhibits myopic behaviors such as oscillatory paths in dynamic environments [13]. As a step forward, optimal reciprocal collision avoidance (ORCA) and social force account for the velocities of agents. ORCA models other agents as velocity obstacles and assumes that agents avoid each other under the reciprocal rule [59, 60]. SF models the interactions between the robot and other agents using attractive and repulsive forces [20]. However, the hyperparameters of the model-based approaches are sensitive to crowd behaviors and thus need to be tuned carefully to ensure good performance [39, 13]. In addition, model-based methods are prone to failures, such as the freezing problem, if the assumptions such as the reciprocal rule are broken [57, 38]. In contrast, while our method also models these interactions, we learn the hyperparameters of the model from trial and error with minimal assumptions on human behaviors posed by model-based methods.
@@ II-B Learning-based methods
Learning-based approaches have been widely used for navigation in dynamic environments to reduce hyperparameter tuning efforts and the number of assumptions. One example is supervised learning from expert demonstrations of desired behaviors, where the expert ranges from model-based policies [40, 56, 68], human teleoperators in simulators [49], to real pedestrians [30, 3]. Supervised learning does not require explorations of the state and action spaces of the robot, yet the performance of learned policy is limited by the quality of expert demonstrations.
Another line of work takes advantage of crowd simulators and learns policies with RL. Through trial and error, RL has the potential to learn robot behaviors that outperform model-based approaches and expert demonstrations [67]. For example, Deep V-Learning first uses supervised learning with ORCA as the expert and then uses RL to learn a value function for path planning [9, 16, 7, 10, 70]. However, Deep V-Learning assumes that state transitions of all agents are known without uncertainty. In addition, since the networks are pre-trained with supervised learning, they share the same disadvantages with OCRA, which are hard to correct by RL [36]. To address these problems, more recent efforts leverage model-free RL without supervised learning or assumptions on state transitions [36, 46, 38, 70]. Since the state transition probability of humans is uncertain, directly learning a policy network without explicitly modeling state transitions is more suitable for navigation [53]. However, the aforementioned RL works have at least one of the following two problems: (1) They focus on navigation in open spaces and isolate agents from static constraints, posing difficulties when deploying robots in the real-world; (2) They ignore all or part of interactions among entities in the scene, which are important for robot navigation in dense crowds and highly constrained environments. We discuss how prior works address the above two problems in Sec. II-C and Sec. II-D respectively.
@@ II-C Scene representation for constrained environments
To deal with both static constraints and humans, some methods such as DWA [17] and DS-RNN [36, 8] use groups of small circles to represent the contours of both static obstacles and humans. While straightforward, groups of circle representation has the following problems: 1. Humans and obstacles are treated in the same way but have different effects on robot decision making. This design choice leads to suboptimal policy as we will show in Sec. V and VI. 2. Since the length of input vectors is proportional to the number of obstacles and humans, the hyperparameters of navigation algorithms easily overfit to a specific environment setting, causing performance degradation when the densities of obstacles and humans change.
Other learning-based approaches use raw sensor images or point clouds to represent the whole environment [48, 15, 49, 71, 53]. These end-to-end (e2e) pipelines have made promising progress in simulation. However, generalizing these e2e methods to real-world scenarios requires high-fidelity simulators [58, 50, 31] and/or datasets [42, 30, 21]. Since accurate pedestrian simulation is technically difficult and expensive [31], learning a deployable e2e policy in dense and interactive crowds remains an open challenge. To circumvent this problem, other works leverage processed state information such as occupancy maps [67, 69]. However, compared to human circles or point clouds, occupancy maps expand the state space, which reduces training efficiency. In addition, the discretization of agent positions limits the ability of occupancy maps to capture slow agent movements. For this reason, we develop a structured input representation, consisting of human circles and processed obstacle point cloud, that only includes the most essential information for navigation. This input representation leads to strong performance in both simulation and real-world while only requiring a low-fidelity simulator.
@@ II-D Graph attention networks for multi-agent interactions
In recent years, attention mechanisms have demonstrated success in various fields [61, 62, 63]. Vaswani et al. propose a transformer with self-attention mechanism that achieves state-of-the-art performance in machine translation [61]. Later, graph attention networks show the effectiveness of attention on learning relationships of data with graphical structures such as citation networks [62]. Inspired by these works, researchers in trajectory prediction and crowd navigation have found that attention networks are also well-suited to capture interactions amongst agents and entities, which contain essential information for multi-agent tasks [63, 26, 7, 36, 19]. For each agent, these works compute attention scores for all neighboring agents. Due to the permutation invariance property, attention scores better capture the importance of pair-wise relationships than encoding the states of all agents with a concatenation operation or an recurrent network [9, 16, 8].
@@
More specifically, a line of works uses a robot-human attention network to determine the relative importance of each human to the robot [7, 32, 36, 34]. However, interactions among humans, which can also influence the robot, are not explicitly modeled. To this end, another line of works use a homogenous graph network to include both RH and HH interactions [6, 70, 25]. However, since these works feed RH and HH features to a single attention network, the resulting robot policy has difficulty reasoning the specificity of each type of feature, which limits the robot’s ability to adapt to different interactions, as demonstrated in [5] and our experiments in Sec. V. To deal with this issue, some recent works use multiple attention networks to encode RH and HH features separately, leading to improved navigation performance in dense and interactive crowds [35, 65, 70]. However, these works only deal with open-world social navigation and thus ignore the interactions between agents and obstacles, which are common in real-world navigation scenarios.
@@
To handle obstacles, Liu et al. [37] and Chen et al. [5] treat the humans, the robot, and static objects as nodes in graph neural networks to model the spatio-temporal interactions among all observed entities. Different from Chen et al. [5] where a robot needs to reach an object such as a chair, the robot in this work only needs to avoid all objects. Thus, Chen et al. [5] represent static obstacles as a semantic map with a class label for each object, whereas we represent all observed static obstacles as a point cloud. In Liu et al. [37], representing each obstacle as a geometric shape requires further map processing, such as obstacle segmentation. Thus, we treat the point cloud consisting of all obstacles as a single node in our st-graph, which reduces both the processing requirement and the size of the graph.
@@ III-B Scene representation
To aid policy learning with the complicated state space involving multiple agents and entities, we develop a structured representation that splits human and obstacle states to account for their different effects on robot decision making. In Fig. 3, at each timestep $t$, we split a scene into a human representation ${h}^{t}_{1},...,{h}^{t}_{n}$, and an obstacle representation ${o}^{t}$. Each human has its own state since it can move independently from one another, whereas all obstacles are mixed together since they are all static. In human representation (Fig. 3(b)), the position and velocity of each human are detected using off-the-shelf human detectors [29, 52, 66]. Tracking each human is not required, further simplifying perception requirements. By representing each human as a low-dimentional state vector, we abstract away detailed information such as gaits and appearance, which are difficult to simulate accurately [58, 45].
@@ IV-A Heterogeneous Spatio-Temporal Graph
The subtle yet highly dynamic interactions among agents and entities are important factors that makes crowd navigation difficult. To model these interactions in a structured manner, we formulate the navigation scenario as a heterogeneous st-graph. In Fig. 4a, at each timestep $t$, our heterogeneous st-graph $\mathcal{G}^{t}=(\mathcal{V}^{t},\mathcal{E}^{t})$ consists of a set of nodes $\mathcal{V}^{t}$ and a set of edges $\mathcal{E}^{t}$. The nodes include the detected humans $h_{1}^{t},...,h_{n}^{t}$ and the robot $w^{t}$. In addition, an obstacle node $o^{t}$ represents the point cloud of all obstacles as a whole. At each timestep $t$, the spatial edges that connect different nodes denote the spatial interactions among nodes. Different spatial interactions have different effects on robot decision-making. Specifically, since we have control of the robot but not the humans, RH interactions have direct effects while HH interactions have indirect effects on the robot actions. As an example of indirect effects, if human A aggressively forces human B to turn toward the robot’s front, the robot has to respond as a result of the interaction between A and B. Additionally, since the agents are dynamic but the obstacles are static, interactions among agents are mutual while the influence of static obstacles on agents is one-way. Thus, we categorize the spatial edges into three types: HH edges (blue in Fig. 4), obstacle-agent (OA) edges (orange), and RH edges (red). The three types of edges allow us to factorize the spatial interactions into HH, OA, and RH functions. Each function is parameterized by a neural network that has learnable parameters. Compared to previous works that ignore some edges [7, 36, 35], our method allows the robot to reason about all observed spatial interactions that exist in constrained and crowded environments.
@@
To reduce the number of parameters, the same type of edges in Fig. 4(a) share the same function parameters. This parameter sharing is important for the scalability of our graph because the number of parameters is kept constant when the number of human changes [28].
@@ IV-B1 Attention among agents
The attention modules assign weights to all edges that connect to a robot or a human node, allowing the node to pay attention to important edges or interactions. The two attention networks are similar to the scaled dot-product attention with a padding mask [61], which computes the attention score using a query $Q$ and a key $K$, and applies the normalized score to a value $V$, which results in a weighted value $v$.
@@
where $c_{1},...,c_{n}$ are constants that reflect the effect of the scaling factor $d$ and the softmax function. Thus, the value of the $n$-th missing human $v_{n}$ is eliminated by the mask $M^{t}$ and thus does not affect the resulting weighted human embedding $v_{HH}^{t}$. The mask that indicates the visibility of each human prevents attention to undetected humans, which is common in crowd navigation due to the partial observability caused by the limited robot sensor range, occlusions, imperfect human detectors, etc [46]. Additionally, the mask provides unbiased gradients to the networks, which stabilizes and accelerates the training [41, 35].